OpenStudy (anonymous):

What are the x-intercepts of the quadratic function? y = x2 – x – 2 A. 2 and 1 B. –2 and 1 C. 2 and –1 D. 1

OpenStudy (anonymous):

@DecentNabeel

OpenStudy (decentnabeel):

put y=o then solve it for x intercept

OpenStudy (lynfran):

x=2 x=-1

OpenStudy (decentnabeel):

y = x2 – x – 2 0 = x2 – x – 2 you can sole it

OpenStudy (anonymous):

B

OpenStudy (decentnabeel):

\[0=x^2-x-2\quad :\quad x=2,\:x=-1\]

OpenStudy (decentnabeel):

C

OpenStudy (anonymous):

that was going to be my 2nd option if i was wrong so thank you very much

OpenStudy (decentnabeel):

no problem :(

OpenStudy (anonymous):

whats with the frowny face?